Circular No. 6732

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970828 T. Murakami and Y. Ueda, Institute of Space and Astronautical Sci-

ence; A. Yoshida and N. Kawai, Institute of Physical and Chemical Research; and F. E. Marshall, R. H. D. Corbet, and T. Takeshima, Goddard Space Flight Center, report on behalf of the ASCA and RXTE teams: “We analyzed all of the ASCA observational data from Aug. 29.91 to 30.85 UT (net exposure 36 000 s). After the attitude determination, we refined the ASCA x-ray position (reported on IAUC 6729) to α = 18^{h}08^{m}32^{s}.2, δ = +59^{o}18′54′′ (equinox 2000.0; error radius 0′.5). Its average flux in the band 2–10 keV is 4 × 10^{−13} erg cm^{−2} s^{−1}. The fluxes clearly show fading behavior. This strongly supports the conclusion that the ASCA x-ray source is a fading x-ray source associated with GRB 970828 (IAUC 6726, 6728). Considering the flux reported by the PCA/RXTE observations (IAUC 6727), together with the ASCA fluxes, the x-ray source appears to be fading consistent with a time power-law having an index of ∼ −1.4.”

GRB 970228 S. R. Kulkarni, S. G. Djorgovski, and J. C. Clemens, California Insti-

tute of Technology, report: “On Aug. 13 UT, we used the Low Resolution Imaging Spectrograph mounted on the Keck II telescope to acquire spectra at the position of the GRB 970228 optical transient reported on IAUC 6584. We placed a 1′′.5 slit (p.a. 87^{o}) at the location specified on IAUC 6588 (Metzger et al.), and used a low-resolution grating (300 lines/mm blazed at 500 nm) to acquire exposures totalling 1800 s. A spectral trail is well detected in this location. The spectrum has some weak features but no strong emission lines. Extraction of a reliable magnitude of the object from the spectrum is difficult; the detected flux is consistent with an object of magnitude R ∼ 24–25.”

S. G. Djorgovski, S. R. Kulkarni, R. R. Gal, and S. C. Odewahn, California Institute of Technology; and D. A. Frail, National Radio Astronomy Observatory, report: “We obtained deep CCD images of the field of GRB 970228 with the Palomar 5-m Hale telescope, in subarcsecond seeing, on Sept. 4 UT. We confirm the existence of a faint object (magnitude R ∼ 25.5) associated with the optical transient (IAUC 6588, 6606) and recently detected spectroscopically with the Keck II 10-m telescope by Kulkarni et al. (see above). The brightness of the source is fully consistent with the previous measurements of the nebulosity adjacent to the optical transient. We interpret this to be the host galaxy of GRB 970228.”

1997 September 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green